OpenStudy (deswilli151):

If y varies directly as x, and y is 18 when x is 5, which expression can be used to find the value of y when x is 11?

OpenStudy (wanell123):

whats the answer choices

OpenStudy (wanell123):

5/11 = 18/y 5 y = 18 · 11 5 y = 198 y = 198 : 5 y = 39.6